OpenStudy (anonymous):

A solution has 30.0 mg K2SO4/mL of solution. What is the molarity of K+ in this?

OpenStudy (anonymous):

What I've done so far: 30 mg/mL = 30,000 mg/L = 30 g / L. K2SO4 = 174.27 g/mol. 30/174.27 = 0.17215 0.17215 x (2 mole K/1 mole K2SO4) = 0.34430M K But this is not correct?

OpenStudy (anonymous):

Wow, nevermind. I just deleted 2 numbers off the end and it worked.

OpenStudy (anonymous):

Good work. I was going to ask if you enter the answer online, which it seems that you are, the computer wants exactly the correct sig figs.

OpenStudy (anonymous):

Yep, I don't see how 3 sig figs is correct, but I had 5 chances to enter an answer, so I messed with the sig figs and it worked.

OpenStudy (anonymous):

30.0 is 3 sig figs. What site is your teacher using?

OpenStudy (anonymous):

Webassign

OpenStudy (anonymous):

So they went straight from 30.0 to the answer?

OpenStudy (anonymous):

When I used web assign for math there was a practice option and I was able to see the answer which helped me figure out the format the computer wanted. Do you have that option?

OpenStudy (anonymous):

Sig figs for multiplication and division (addition and subtraction is different) is based on the lowest sig figs. So your answer needs to be rounded to the same digits as in 30.0

OpenStudy (anonymous):

Yea, I do use the lowest, but in my work I had 0.17215 x (2/1) as a step of its own, so I said 5 sig figs.

OpenStudy (anonymous):

I guess they want it all in 1 step.

OpenStudy (anonymous):

Are you free for a little while longer? Could use some help understanding some of the other 4 I'm stuck on.

OpenStudy (anonymous):

If let's say one has 6 figs and one has 4 sig figs your final answer should have 4 sig figs. The accuracy or precision I'm not sure which term is correct can only be as precise accurate as the lowest one. I can help with one more and then need to call it a night.

OpenStudy (anonymous):

Oh, ok, I thought it was per equation.

OpenStudy (anonymous):

And thanks, I'll post one of the others then.